Enlarge existing rectangle such that total area is twice as large. Current rectangle dimensions are 16 feet by 25 feet. Add 11 feet more to the long side than to the short side. Find new dimensions.
Answer by stanbon(75887)

Enlarge existing rectangle such that total area is twice as large. Current rectangle dimensions are 16 feet by 25 feet. Add 11 feet more to the long side than to the short side. Find new dimensions.
====
Original Area: 16*25 = 400 sq ft.
====
New dimensions:
25+11 + x = 36+x
16+x
---
Solve for "x":
(36+x)(16+x) = 800
x^2+52x+576 = 800
x^2+52x-224 = 0
(Positive solution:
x = 4 ft
---
New dimensions:
25+11 + 4 = 40 ft
16+4 = 20 ft
===================
New Area: 20*40 = 800 sq ft
================================
Cheers,
Stan H.
---------------------